Solve the Initial Value problem
y"− 4y′ + 3y = 0,y( 0 )= 7,y′ (0) = 11
1
Expert's answer
2021-07-23T07:49:36-0400
"y''-4y'+3y=0"

The characteristic equation is


"r^2-4r+3=0"

"(r-1)(r-3)=0"

"r_1=1, r_2=3"

The general solution of the homogeneous differential equation is


"y(t)=c_1e^t+c_2e^{3t}"

"y(0)=7: 7=c_1e^0+c_2e^{3(0)}"

"c_1+c_2=7"

"y'(t)=c_1e^t+3c_2e^{3t}"

"y'(0)=11: 11=c_1e^0+3c_2e^{3(0)}"


"c_1+3c_2=11"

Then


"c_1=5, c_2=2"

Solution of the Initial Value problem is


"y(t)=5e^t+2e^{3t}"